# Trajectory Class Fluctuation Theorem

### v2

@@ Abstract
The Trajectory Class Fluctuation Theorem (TCFT) substantially strengthens the Second Law of Thermodynamics—that, in point of fact, can be a rather weak bound on resource fluxes. Practically, it improves empirical estimates of free energies, a task known to be statistically challenging and has diagnosed successful and failed information processing in experimentally-implemented Josephson-junction information engines. The development here justifies that empirical analysis, explicating its mathematical foundations.
@@ II.1 Model, Probability Densities, and Time Reversal
We require that the net work $W({\overrightarrow{z}})$ is a function of trajectory. This can be achieved if $T$ and ${\mathcal{Z}}$ are sufficiently fine. (For example, $T=[0,\tau]$ and ${\mathcal{Z}}$ is the system’s set of microstates.) Then, by conservation of energy, heat $Q({\overrightarrow{z}})$ must also be a function of trajectory and we have the First Law for each trajectory:
@@
where $F^{\text{eq}}_{t}$ is the equilibrium free energy for the system at time $t$. Note that if ${\mathcal{Z}}$ is discrete, then a state probability density evaluated at a state is in fact the corresponding probability of that state. That is, integration over discrete spaces can simply be taken to be summation.
@@
A solution can be found with measure theory but we treat the subject only briefly. A measure is a function on the regions of a space that returns an amount of some “quantity”, such as probability, in a region. We choose a particular measure on trajectory space and refer to it as a base measure. A probability density is then a function that, when Lebesgue-integrated via the base measure over a region of space, yields the corresponding probability of that region. Defining appropriate base measures on a continuous-time trajectory space is rather technical, though, and so we leave the discussion to a sequel.
@@ III.1 Trajectory Classes
Every trajectory class is a subset of ${\overrightarrow{\mathcal{Z}}}$ for which the forward and reverse processes assign probability. Formally, the set of all trajectory classes $\mathcal{C}$ must constitute a $\sigma$-algebra over the trajectories ${\overrightarrow{\mathcal{Z}}}$. However, to stay with the physics, the following does not focus on measure-theoretic details. (A sequel focuses on the latter.) Instead, we first discuss several example classes and thereafter interpret any subset of ${\overrightarrow{\mathcal{Z}}}$ of practical interest to be part of the assumed $\sigma$-algebra and, therefore, to be a trajectory class.
@@
The singleton trajectory classes of Example (6) provides one instance where the model of the system determines whether a trajectory class exists. For a finite number of times $T$, we can assume the singleton trajectory classes exists. However, for technical reasons, such trajectory classes often fail to exist for continuous-time processes. See Sec. VII for more examples as they apply to known results.
@@ III.2 Trajectory Class Quantities
Experimentally, one might repeatedly conduct the reverse experiment to estimate the physical reverse process probability ${R^{\prime}}({C^{\dagger}})$ of obtaining a trajectory in the class ${C^{\dagger}}$, where:
${C^{\dagger}}$ =$=\{{{\overrightarrow{z}}^{\dagger}}|{\overrightarrow{z}}\in C\}~{},$
is the reverse of class $C$. An empirical estimate for the formal reverse-process probability of $C$ is then given by the following identity:
@@ III.5 Scope: Zero Probability Classes
As $\epsilon$ decreases, the work distribution for $C^{\prime}$ necessarily narrows, so that $e^{-\beta(W-\Delta F^{\text{eq}})}$ approaches $e^{-\beta(\widetilde{W}-\Delta F^{\text{eq}})}$. And, if the work distributions for the forward and reverse processes are smooth, then ${R}(C^{\prime})$ and ${P}(C^{\prime})$ will eventually shrink at the same, constant rate. In this case, ${R}(C^{\prime})/{P}(C^{\prime})$ converges to a ratio of work densities at $\widetilde{W}$.
@@
In fact, the TCFT introduced here can be strengthened to directly address classes of zero probability without the need for approximations or limiting schemes. This strengthening is done with the measure-theoretical notion of conditional expectation. However, an exposition requires a more thorough treatment of measure theory and so we treat it in a sequel.
@@ IV.1 Trajectory Class Second Law
Thus, the class reverse surprisal  ${{{\Theta}_{C}}}$—a quantity that only depends on information about trajectories coarse-grained to the class of interest—bounds the class averaged entropy difference ${{\left\langle{{\Sigma}}\right\rangle_{C}}}$.
@@ V.1 Constant Free-Energy Differences
This equality relates free-energy differences to statistics on works and class probabilities. Jarzynski’s Equality Eq. (15) is the special case where $C={\overrightarrow{\mathcal{Z}}}$, ${\rho}={\pi}_{0}$ and ${\sigma}={\pi}_{\tau}$.
@@ VI Statistical Freedom from the Tyranny of the Rare
When estimating statistical quantities from data, rare events can dominate sample averages [12, 13]. Consider the following case in point. By empirically estimating the exponential average work ${\left\langle{e^{-\beta W}}\right\rangle_{{\overrightarrow{\mathcal{Z}}}}}$ for a thermodynamic transformation, one can estimate the equilibrium free energy difference $\Delta F^{\text{eq}}$ via Eq. (15)’s Jarzynski’s Equality. However, this can require thorough sampling of very rare events [12]. The ECFT  of Eq. (2) can aid in solving this statistical challenge by removing consideration of these rare but work-dominant trajectories.
@@ VII Related Results
We now turn to the burgeoning collection of already-established fluctuation theorems. As noted, many existing fluctuation theorems are special cases of the TCFT.
@@ VIII Discussion
The preceding provided guidance when selecting appropriate trajectory classes for a given process of interest. To achieve a much stronger bound on entropy difference than the traditional ensemble-average Second Law, Sec. IV.3 proposed choosing a partition of all trajectories into classes that resulted in narrow entropy difference distributions for each class. And, to avoid dominating rare events when calculating free-energy differences (Sec. VI), we recommended classes that are common in both the forward and reverse processes and that have narrow entropy-difference distributions. That said, the most effective classes for these tasks appear to be specific to the particular processes of interest. Procedures to identify these classes for arbitrary processes remain an open problem.
@@ Appendix C Free Energy Estimate Distribution
If we wish to estimate the change in free energy from the work distributions of a collection forward and reverse experiments that start in equilibrium, the TCFT provides a relation for the free-energy differences of each trajectory class $C$:
@@
To evaluate the free-energy difference estimate, we note $\Delta\tilde{f}_{C}$ is itself a function of the estimated probability of realizations of your experiment that start in $A$:

### v3

@@ Abstract
The Trajectory Class Fluctuation Theorem (TCFT) substantially strengthens the Second Law of Thermodynamics—that, in point of fact, can be a rather weak bound on resource fluxes. Practically, it improves empirical estimates of free energies, a task known to be statistically challenging, and has connected the microscopic dynamics with the mesoscopic information processing in experimentally-implemented Josephson-junction information engines. The development here justifies that empirical analysis, explicating its mathematical foundations.
@@ II.1 Model, Probability Densities, and Time Reversal
We require that the net work $W({\overrightarrow{z}})$ is a function of trajectory. This can be achieved if $T$ and ${\mathcal{Z}}$ are sufficiently refined. (For example, $T=[0,\tau]$ and ${\mathcal{Z}}$ is the system’s set of microstates.) Then, by conservation of energy, heat $Q({\overrightarrow{z}})$ must also be a function of trajectory and we have the First Law for each trajectory:
@@
where $F^{\text{eq}}_{t}$ is the system’s equilibrium free energy at time $t$ and $\beta$ is the inverse temperature as denoted in statistical mechanics. Note that if ${\mathcal{Z}}$ is discrete, then a state probability density evaluated at a state is in fact the corresponding probability of that state. That is, integration over discrete spaces can simply be taken to be summation.
@@
A solution can be found via measure theory but we treat the subject only briefly here. (A sequel provides the details [18].) A measure is a function on the regions of a space that returns an amount of some “quantity”, such as probability, in a region. We choose a particular measure on trajectory space and refer to it as a base measure. A probability density is then a function that, when Lebesgue-integrated via the base measure over a region of space, yields the corresponding probability of that region. Defining appropriate base measures on a continuous-time trajectory space is rather technical, though, and so we leave the discussion to the sequel.
@@ III.1 Trajectory Classes
Every trajectory class is a subset of ${\overrightarrow{\mathcal{Z}}}$ for which the forward and reverse processes assign probability. Formally, the set of all trajectory classes $\mathcal{C}$ must constitute a $\sigma$-algebra over the trajectories ${\overrightarrow{\mathcal{Z}}}$. However, to stay with the physics, the following does not focus on measure-theoretic details. (The sequel focuses on the latter.) Instead, we first discuss several example classes and thereafter interpret any subset of ${\overrightarrow{\mathcal{Z}}}$ of practical interest to be part of the assumed $\sigma$-algebra and, therefore, to be a trajectory class.
@@
The singleton trajectory classes of Example (6) provides one instance where the model of the system determines whether a trajectory class exists. For a finite number of times $T$, we can assume the singleton trajectory classes exist. However, for technical reasons, such trajectory classes often fail to exist for continuous-time processes. See Sec. VII for more examples as they apply to known results.
@@ III.2 Trajectory Class Quantities
The reverse of a trajectory class $C$ is defined as:
${C^{\dagger}}$ =$=\{{{\overrightarrow{z}}^{\dagger}}|{\overrightarrow{z}}\in C\}~{}.$
Then the physical reverse probability ${R^{\prime}}({C^{\dagger}})$ of obtaining trajectory class ${C^{\dagger}}$ during the reverse experiment is equal to the formal reverse probability ${R}(C)$:
@@
With the above equality, we can then obtain an empirical estimate of ${R}(C)$ from reverse experiment data.
@@ III.5 Scope: Zero Probability Classes
As $\epsilon$ decreases, the work distribution for $C^{\prime}$ necessarily narrows, so that $e^{-\beta(W-\Delta F^{\text{eq}})}$ approaches $e^{-\beta(\widetilde{W}-\Delta F^{\text{eq}})}$. And, if the work distributions for the forward and reverse processes are continuous functions of work value, then ${R}(C^{\prime})$ and ${P}(C^{\prime})$ will eventually shrink at the same, constant rate. In this case, ${R}(C^{\prime})/{P}(C^{\prime})$ converges to a ratio of work densities at $\widetilde{W}$.
@@
In fact, the TCFT introduced here can be strengthened to directly address classes of zero probability without the need for approximations or limiting schemes. This strengthening is done with the measure-theoretical notion of conditional expectation. However, an exposition requires a more thorough treatment of measure theory and so we treat it in the sequel.
@@ IV.1 Trajectory Class Second Law
Thus, the class reverse surprisal  ${{{\Theta}_{C}}}$—a quantity that only depends on ${P}(C)$ and ${R}(C)$— bounds the class averaged entropy difference ${{\left\langle{{\Sigma}}\right\rangle_{C}}}$.
@@ V.1 Constant Free-Energy Differences
This equality relates free-energy differences to statistics on works and class probabilities. Jarzynski’s Equality Eq. (15) is the special case where $C={\overrightarrow{\mathcal{Z}}}$, ${\rho}={\pi}_{0}$, and ${\sigma}={\pi}_{\tau}$.
@@ VI Statistical Freedom from the Tyranny of the Rare
When estimating statistical quantities from data, rare events can dominate sample averages [12, 13]. This can be particularly problematic when the events are associated with large resources. Consider the following case in point. By empirically estimating the exponential average work ${\left\langle{e^{-\beta W}}\right\rangle_{{\overrightarrow{\mathcal{Z}}}}}$ for a thermodynamic transformation, one can estimate the equilibrium free energy difference $\Delta F^{\text{eq}}$ via Eq. (15)’s Jarzynski’s Equality. However, this can require thorough sampling of very rare events [12]. The ECFT  of Eq. (2) can aid in solving this statistical challenge by removing consideration of these rare but work-dominant trajectories.
@@ VII Related Results
We now turn to the burgeoning collection of previously established fluctuation theorems. As noted, many existing fluctuation theorems are special cases of the TCFT.
@@ VIII Discussion
The preceding provided guidance when selecting appropriate trajectory classes for a given process of interest. To achieve a much stronger bound on entropy difference than the traditional ensemble-average Second Law, Sec. IV.3 proposed choosing a partition of all trajectories into classes that resulted in narrow entropy difference distributions for each class. And, to avoid dominating rare events when calculating free-energy differences (Sec. VI), we recommended classes that are common in both the forward and reverse processes and that have narrow entropy-difference distributions. That said, the most effective classes for these tasks appear to be specific to the particular processes of interest. Developing procedures to identify these classes for arbitrary processes remain an open problem.
@@ Appendix C Free Energy Estimate Distribution
If we wish to estimate the change in free energy from the work distributions of a collection of forward and reverse experiments that start in equilibrium, the TCFT provides a relation for the free-energy differences of each trajectory class $C$:
@@
To evaluate the free-energy difference estimate, we note $\Delta\tilde{f}_{C}$ is itself a function of the estimated probability of realizations of the experiment that starts in $A$: